If the relative error of m is 1% and of h is .5%, what is the error of log(mh)?
Answer #1

log(mh) = log(m) + log(h)

diffrentiating both side

d(mh)/mh = dm/m + dh/h

d(mh)/mh = 0.01 + 0.005 = 0.015

=> error of log(mh) = 1.5%